# Collider imprint of vector-like leptons in light of anomalous magnetic moment and neutrino data

### v1

@@ 3 Scan Results
After conducting a random scan across the parameter space of the model as defined in Table 3, we have identified 328 solutions that satisfy all bounds listed in Table 2 along with restrictions on VLL masses and milder constraints originating from the anomalous magnetic moments of the muon and the electron to orient the scan towards the physics case studied. Many of these solutions hence exhibit significant deviations from the SM predictions for the anomalous magnetic moments of both the muon and the electron, potentially offering a pathway to reconciling theory with data. Specifically, these solutions feature $\Delta{a}_{\mu}$ values reaching up to $3\sigma$, as defined by Eq. (8), and $\Delta{a}^{\rm Rb}_{e}$ values showing positive deviations of up to $2\sigma$, aligning with the results obtained from measurements carried out from Rubidium atoms and presented in Eq. (9). However, none of the obtained solutions predicts a negative deviation for $a_{e}$, demonstrating that our model cannot account for the value of $a_{e}$ derived from measurements carried out on Cesium atoms. In the following, we investigate these findings in detail, first relative to the anomalous magnetic moments of the muon and the electron in section 3.1, and next in the context of neutrino data in section 3.2. Finally, we verify in section 3.3 that adding multiplets of VLLs to the theory does not impact electroweak precision tests.
@@ 3.1 Anomalous magnetic moments of the muon and the electron
To further explore the impact of all model’s free parameters on $\Delta a_{\mu}$ and $\Delta a_{e}$, we now examine how they correlate with predictions for the value of the electron and muon anomalous magnetic moments. Figure 4 shows heat maps based on the pairwise Pearson correlation coefficients JamesISL ; Lista:2017jsy between these parameters, and $\Delta a_{e}$ (left) and $\Delta a_{\mu}$ (right). Pearson correlation coefficients measure the linear relationship between two quantities, by calculating the ratio of their covariance to the product of their standard deviations. According to this definition, the coefficient lies within the range $[-1,1]$, where a value equal to $1$ suggests a perfect positive linear relationship (the two quantities exhibiting a similar behaviour), and $-1$ a perfect negative linear relationship (the two quantities exhibiting an opposite behaviour), while 00 represents no linear relationship. However, such coefficients ignore other types of (non-linear) relationships between quantities. The corresponding formula for calculating the Pearson coefficients between two parameters $x$ and $y$ across a given dataset is given by
@@ 3.2 Properties of the neutrino sector
The introduction of VLLs has implications for the neutrino sector, providing a mechanism to generating neutrino masses. While compelling evidence from oscillation experiments indicates that their mass is non-zero SNO:2002tuh ; Super-Kamiokande:2005mbp , the actual mass spectrum is unresolved. It can nevertheless be constrained thanks to cosmology. Indeed, the abundance of cosmic neutrinos and their contribution to the universe’s total energy density during its early stages enables the study of their characteristics. Specifically, the sum of the neutrino masses $\sum m_{\nu}$ influences the cosmic microwave background, supernovæ features, large scale structure, and big bang nucleosynthesis Loureiro:2018pdz . Present data has yielded an approximate upper limit of $\sum m_{\nu}<0.26$ eV, a constraint that we have incorporated in our parameter space scans (see section 2). In our analysis of the neutrino sector, we have considered scenarios with the normal ordering (NO) and inverted ordering (IO) of the Majorana neutrino masses, and we have explored all mixing possibilities of the SM neutrinos with the neutral components of the VLL doublets and neutral singlets. The constraints imposed on the neutrino sector during the scan are detailed in Table 4.
Cosmological constraints Neutrino masses (IO) Neutrino masses (NO)
$\Sigma\,m_{\nu}\,\leq\,$0.26 eV $m_{\nu_{1}}\in[8.37,8.85]\times 10^{-3}$ eV $m_{\nu_{2}}\in[8.46,8.88]\times 10^{-3}$ eV
$m_{\nu_{2}}\in[4.93,5.04]\times 10^{-2}$ eV $m_{\nu_{3}}\in[4.88,5.02]\times 10^{-2}$ eV
@@
In order to assess which model parameters have the most significant impact on the neutrino sector, we present a heat map in the left panel of figure 6 based on the pairwise Pearson correlation coefficients between the mass of the neutrino $\nu_{3}$ (chosen for illustrative purposes) and various model parameters relevant for neutrino masses at one loop. It turns out that the parameter $\mathbf{Y}_{\mathbf{\nu}}^{33}$, governing the Yukawa interactions between ordinary leptons and the right-handed neutrino field, has the largest effect. Turning our attention to the value of the one-loop corrections $\Delta m_{\nu_{3}}$ to the tree-level mass, we observe, in the right panel of the figure, that the value of $\mathbf{Y}_{\mathbf{\nu}}^{33}$ plays a minimal role on the neutrino mass for $\mathbf{Y}_{\mathbf{\nu}}^{33}\lesssim 10^{-12}$. However, for larger values, these effects become more pronounced as one-loop contributions $\Delta m_{\nu_{3}}$ must compensate for a too large tree-level mass. This easily leads to predictions that conflict with neutrino oscillation data, as visible from the figure. Therefore, achieving agreement with experimental constraints necessitates a very weak interaction term between the right-handed neutrino field $\nu$ and the SM-like lepton doublet $L$.
@@ 4 Collider phenomenology
In this section, our focus is on analysing the phenomenological implications at high-energy hadronic colliders of scenarios that meet all previously imposed constraints. We examine two representative benchmark setups and demonstrate that they yield discernible signals amidst the overwhelming background of the SM. For the first scenario (see section 4.1), we exploit a signature characterised by the production of a significant number of charged leptons, while for the second one (see section 4.2), we leverage the production of substantial missing energy. In general, studying these two signals is sufficient to obtain excellent collider handles on model scenarios consistent with all constraints examined in section 3.
[…]
@@ 4.1 A signature involving multiple charged leptons
$p\,\,p\ \ \rightarrow\ \ e^{-}_{9}\,\,e^{+}_{9}+X\ \ \rightarrow\ \ \mu^{-}\,Z\,\mu^{+}\,\,Z+X\ \ (\rightarrow\mu^{-}\ell^{+}\ell^{-}\,\mu^{-}\ell^{\prime+}\ell^{\prime-}+X)\,.$ (11)
@@
$p\,\,p\ \rightarrow\ h\,\,Z+X\,,\quad p\,\,p\ \rightarrow\ Z\,\,Z\,\,Z+X\quad\text{and}\quad p\,\,p\ \rightarrow\ t\,\,\bar{t}\,\,Z\,\,Z+X\,,$ (12)
which all lead to a six-lepton final state once $Z$-boson and Higgs-boson leptonic decays are accounted for. The generated events are re-weighted to ensure that the corresponding hard-scattering cross sections, including the branching ratios related to leptonic Higgs and $Z$ boson decays, are accurate to next-to-next-to-leading order (NNLO) for the first process Gauld:2021ule , and NLO for the last two processes Alwall:2014hca . This yields cross section values of $\sigma_{Zh}=0.455~{}\text{ab}$, $\sigma_{ZZZ}=0.0602~{}\text{fb}$ and $\sigma_{t\bar{t}ZZ}=0.371~{}\text{ab}$, that we determine by multiplying the leading-order predictions returned by MadGraph5_aMC@NLO by constant $K$-factors extracted from the literature.
@@
$n_{s}=28.43\qquad\text{and}\qquad n_{b}=9.24\quad\text{with}\ \ n_{Zh}\simeq 0,\ \ n_{ZZZ}=6.23\ \ \text{and}\ \ n_{t\bar{t}ZZ}=2.99\,.$ (13)
@@
$\begin{split}s=&\ \frac{n_{s}}{\sqrt{n_{b}+\Delta_{b}^{2}}}\,,\\ Z_{A}=&\ \sqrt{2\left\{(n_{s}+n_{b})\ln\left[\frac{(n_{s}+n_{b})(n_{s}+\Delta_{b}^{2})}{n_{b}^{2}+(n_{s}+n_{b})\Delta_{b}^{2}}\right]-\frac{n_{b}^{2}}{\Delta_{b}^{2}}\ln\left[1+\frac{\Delta_{b}^{2}n_{s}}{n_{b}(n_{b}+\Delta_{b}^{2})}\right]\right\}}\ ,\end{split}$ (14)
@@ 4.2 A signature with significant missing transverse energy
$p\,\,p\ \ \rightarrow\ \ e^{-}_{7}\,\,e^{+}_{7}+X\ \ \rightarrow\ \ \tau^{-}\,Z\,\tau^{+}\,\,Z+X\ \ (\rightarrow\tau^{-}\ell^{+}\ell^{-}\,\tau^{-}\ell^{\prime+}\ell^{\prime-}+X)\,.$ (16)
@@
$n_{s}=53.3\qquad\text{and}\qquad n_{b}=23.8\quad\text{with}\ \ n_{Zh}\simeq 0,\ \ n_{ZZZ}=2.3\ \ \text{and}\ \ n_{t\bar{t}ZZ}=21.5\,,$ (17)
@@ 5 Summary and conclusions
In the absence of direct signals of new physics at colliders, indirect constraints play a crucial role in constructing relevant and motivated BSM theories. This often involves reconciling discrepancies between SM predictions and experimental data, such as those observed in the anomalous magnetic moments of the muon and electron that can be further linked to neutrino physics. Our study exploits precisely this. We examined a simplified extension of the SM to which we added three generations of vector-like leptons, each associated with a specific SM family.
After exhaustively listing anomaly-free representations for VLLs and their interaction Lagrangian, we identified parameter space regions compatible with neutrino data and consistent with measurements of the anomalous magnetic moment of the electron and the muon. We determined that this requires, minimally, the inclusion of two weak VLL singlets (one with the quantum numbers of a right-handed electron and the other with the quantum numbers of a right-handed neutrino), along with a weak VLL doublet with hypercharge $-1/2$. The compatible parameter space regions imply distinct collider signatures, prominently featuring six hard leptons accompanied by either significant total transverse energy or missing transverse energy. We performed a detailed phenomenological analysis to demonstrate the sensitivity of future proton-proton colliders operating at a centre-of-mass energy of 100 TeV to these signatures, focusing on two representative benchmark scenarios with VLL masses in the range of $100-400$ GeV. Our findings underscore that if new physics involves VLLs with properties explaining neutrino data and reconciling discrepancies in anomalous magnetic moments of the light-generation leptons, it will yield signals that will be unmistakable at future proton-proton colliders currently under discussion within the high-energy physics community.

### v2

@@ 3 Scan Results
After conducting a random scan across the parameter space of the model as defined in Table 3, we have identified 328 solutions that satisfy all bounds listed in Table 2 along with restrictions on VLL masses and milder constraints originating from the anomalous magnetic moments of the muon and the electron to orient the scan towards the physics case studied. Many of these solutions hence exhibit significant deviations from the SM predictions for the anomalous magnetic moments of both the muon and the electron, potentially offering a pathway to reconciling theory with data. Specifically, these solutions feature $\Delta{a}_{\mu}$ values reaching up to $3\sigma$, as defined by Eq. (16), and $\Delta{a}^{\rm Rb}_{e}$ values showing positive deviations of up to $2\sigma$, aligning with the results obtained from measurements carried out from Rubidium atoms and presented in Eq. (17). However, none of the obtained solutions predicts a negative deviation for $a_{e}$ such that our model cannot explain $a_{e}$ values as extracted from measurements of properties of Cesium atoms measurements. In the following, we investigate these findings in detail, first relative to the anomalous magnetic moments of the muon and the electron in section 3.1, and next in the context of neutrino data in section 3.2. Finally, we verify in section 3.3 that adding multiplets of VLLs to the theory does not impact electroweak precision tests.
@@ 3.1 Anomalous magnetic moments of the muon and the electron
To further explore the impact of all model’s free parameters on $\Delta a_{\mu}$ and $\Delta a_{e}$, we now examine how they correlate with predictions for the value of the electron and muon anomalous magnetic moments. Figure 4 shows heat maps based on the pairwise Pearson correlation coefficients JamesISL ; Lista:2017jsy between these parameters, and $\Delta a_{e}$ (left) and $\Delta a_{\mu}$ (right). Each cell in these heat maps includes the corresponding correlation values (between $-1$ and $1$), and is colored accordingly. The correlations are obtained from a sample containing the ensemble of the 328 scenarios identified by our scanning procedure, each scenario contributing as one input vector that includes the value of each independent parameter of the model (e.g., $\mathbf{\hat{Y}}_{\nu}$, $\mathbf{\hat{Y}}_{\mathcal{S}_{1}}$, etc.) along with those of the observables $\Delta a_{\mu}$ and $\Delta a_{e}$ calculated at one loop by SPheno. In practice, Pearson correlation coefficients are computed with a homemade Python program using a specific NumPy harris2020array array of 328 elements for each component of the input vector, the data being collected within a pandas reback2020pandas ; mckinney-proc-scipy-2010 data frame. Then we use StandardScalar class within the preprocessing package of the sklearn library scikit-learn to standardize whole the dataset, implying that large values of the input parameters can be compared to small values within a correlation calculation. Employing the standardized dataset, we finally proceed with calculating the correlations by applying the corr method to the collected data frame, and generate the plots with the heatmap function of the seaborn library Waskom2021 .
To elaborate on the employed correlation formula, we remind that Pearson correlation coefficients measure the linear relationship between two quantities by calculating the ratio of their covariance to the product of their standard deviations. According to this definition, the coefficient lies within the range $[-1,1]$, where a value equal to $1$ suggests a perfect positive linear relationship (the two quantities exhibiting a similar behaviour), and $-1$ a perfect negative linear relationship (the two quantities exhibiting an opposite behaviour), while 00 represents no linear relationship. Thus the heat maps in figure 4 indicate the linear relationships of $\Delta a_{\mu}$ and $\Delta a_{e}$ with the mass and Yukawa terms of the Lagrangian. However, such coefficients ignore other types of (non-linear) relationships between quantities. The corresponding formula for calculating the Pearson coefficients between two parameters $x$ and $y$ across a given dataset is given by
@@ 3.2 Properties of the neutrino sector
The introduction of VLLs has implications for the neutrino sector, as it can impact the mechanism to generating neutrino masses. We must therefore ensure that the model is compatible with neutrino data. While compelling evidence from oscillation experiments indicates that their mass is non-zero SNO:2002tuh ; Super-Kamiokande:2005mbp , the actual mass spectrum is unresolved. It can nevertheless be constrained thanks to cosmology. Indeed, the abundance of cosmic neutrinos and their contribution to the universe’s total energy density during its early stages enables the study of their characteristics. Specifically, the sum of the neutrino masses $\sum m_{\nu}$ influences the cosmic microwave background, supernovæ features, large scale structure, and big bang nucleosynthesis Loureiro:2018pdz . Present data has yielded an approximate upper limit of $\sum m_{\nu}<0.26$ eV, a constraint that we have incorporated in our parameter space scans (see section 2). In our analysis of the neutrino sector, we have considered scenarios with the normal ordering (NO) and inverted ordering (IO) of the Majorana neutrino masses, and we have explored all mixing possibilities of the SM neutrinos with the neutral components of the VLL doublets and neutral singlets. The constraints imposed on the neutrino sector during the scan are detailed in Table 4.
Cosmological constraints
$\Sigma\,m_{\nu}\,\leq\,$0.26 eV
Neutrino masses (IO)
$m_{\nu_{1}}\in[8.37,8.85]\times 10^{-3}$ eV
$m_{\nu_{2}}\in[4.93,5.04]\times 10^{-2}$ eV
Neutrino masses (NO)
$m_{\nu_{2}}\in[8.46,8.88]\times 10^{-3}$ eV
$m_{\nu_{3}}\in[4.88,5.02]\times 10^{-2}$ eV
@@
In order to assess which model parameters have the most significant impact on the neutrino sector, we present a heat map in the left panel of figure 6 based on the pairwise Pearson correlation coefficients between the mass of the neutrino $\nu_{3}$ (chosen for illustrative purposes) and various model parameters relevant for neutrino masses at one loop. It turns out that the parameter $\mathbf{Y}_{\mathbf{\nu}}^{33}$, governing the Yukawa interactions between ordinary leptons and the right-handed neutrino field, has the largest effect. Turning our attention to the value of the one-loop corrections $\Delta m_{\nu_{3}}$ to the tree-level mass, we observe, in the right panel of the figure, that the value of $\mathbf{Y}_{\mathbf{\nu}}^{33}$ plays a minimal role on the neutrino mass for $\mathbf{Y}_{\mathbf{\nu}}^{33}\lesssim 10^{-12}$. However, for larger values, these effects become more pronounced as one-loop contributions $\Delta m_{\nu_{3}}$ must compensate for a too large tree-level mass. This easily leads to predictions that conflict with neutrino oscillation data, as visible from the figure. Therefore, achieving agreement with experimental constraints necessitates a very weak interaction term between the right-handed neutrino field $\nu$ and the SM-like lepton doublet $L$. Furthermore, our results demonstrate that while neutrino data can be accommodated in a satisfactory manner, the additional degrees of freedom provided by the VLL sector extend the allowed range for the neutrino Yukawa couplings. This could pave the way to a theoretically-acceptable solution to the neutrino mass problematic with larger Yukawa coupling values. The tiny values of the Yukawa couplings indeed demonstrate by themselves that right-handed neutrinos alone cannot be a reason for explaining the neutrino data, as such small values are theoretically not elegant.
@@ 4 Collider phenomenology
In this section, our focus is on analysing the phenomenological implications at high-energy hadronic colliders of scenarios that meet all previously imposed constraints. We examine two representative benchmark setups and demonstrate that they yield discernible signals amidst the overwhelming background of the SM Bissmann:2020lge . For the first scenario (see section 4.1), we exploit a signature characterised by the production of a significant number of charged leptons, while for the second one (see section 4.2), we leverage the production of substantial missing energy. In general, studying these two signals is sufficient to obtain excellent collider handles on model scenarios consistent with all constraints examined in section 3.
To conduct our study, we generate a UFO model Darme:2023jdn with the SARAH package (version 4.15.1) Staub:2013tta ; Goodsell:2017pdq , starting from the model implementation discussed in section 2. This allows us to utilise MadGraph5_aMC@NLO Alwall:2014hca (version 3.2.0) to generate both signal and background events relevant to a potential proton-proton collider operating at a centre-of-mass energy of 100 TeV, and aiming to collect an integrated luminosity of 3 ab-1. We obtain hard-scattering events by convoluting leading-order matrix elements with the next-to-leading-order set of NNPDF 2.3 parton densities Ball:2012cx . Subsequently, these events are matched with parton showering using Pythia 8.3 Bierlich:2022pfr , which we also use to simulate hadronisation. Our simulations match matrix elements including up to one additional parton at the Born level, using the MLM matching scheme Mangano:2006rw ; Alwall:2008qv , and we re-weight events to incorporate higher-order corrections in QCD into the relevant total rates Alwall:2014hca ; Gauld:2021ule ; AH:2023hft . Finally, hadron-level events are processed through the fast detector simulator SFS Araz:2020lnp equipping the MadAnalysis 5 package Conte:2012fm ; Conte:2014zja ; Conte:2018vmg (version 1.10.12). This internally relies on FastJet Cacciari:2011ma (version 3.3.3) for event reconstruction, together with its implementation of the anti-$k_{T}$ jet algorithm Cacciari:2008gp with a radius parameter $R=0.4$. To emulate realistic detector resolution and particle identification, we employ the ATLAS configuration card provided with the SFS module of MadAnalysis 5, which we also use for implementing our collider analysis.
@@ 4.1 A signature involving multiple charged leptons
$\begin{split}p\,\,p\ \ \rightarrow&\ \ e^{-}_{9}\,\,e^{+}_{9}+X\ \ \rightarrow\ \ \mu^{-}\,Z\,\mu^{+}\,\,Z+X\\ &\quad\ \ \rightarrow\mu^{-}\ell^{+}\ell^{-}\,\mu^{-}\ell^{\prime+}\ell^{\prime-}+X\,.\end{split}$ (19)
@@
$\begin{split}&p\,\,p\ \rightarrow\ h\,\,Z+X\,,\quad\\ &p\,\,p\ \rightarrow\ Z\,\,Z\,\,Z+X\,,\quad\\ &p\,\,p\ \rightarrow\ t\,\,\bar{t}\,\,Z\,\,Z+X\,,\end{split}$ (20)
which all lead to a six-lepton final state once $Z$-boson and Higgs-boson leptonic decays are accounted for. The generated events are re-weighted to ensure that the corresponding hard-scattering cross sections, including the branching ratios related to leptonic Higgs and $Z$ boson decays, are accurate to next-to-next-to-leading order (NNLO) for the first process Gauld:2021ule , and NLO for the last two processes Alwall:2014hca . This yields cross section values of $\sigma_{Zh}=0.455\leavevmode\nobreak\ \text{ab}$, $\sigma_{ZZZ}=0.0602\leavevmode\nobreak\ \text{fb}$ and $\sigma_{t\bar{t}ZZ}=0.371\leavevmode\nobreak\ \text{ab}$, that we determine by multiplying the leading-order predictions returned by MadGraph5_aMC@NLO by constant $K$-factors extracted from the literature.
@@
$n_{s}=28.43\qquad\text{and}\qquad n_{b}=9.24$ (21)
with
$n_{Zh}\simeq 0,\ \ n_{ZZZ}=6.23\ \ \text{and}\ \ n_{t\bar{t}ZZ}=2.99\,.$ (22)
@@
$s=\frac{n_{s}}{\sqrt{n_{b}+\Delta_{b}^{2}}}\,,$ (23)
and
$\begin{split}Z_{A}=&\ \Bigg{[}2\bigg{\{}(n_{s}+n_{b})\ln\frac{(n_{s}+n_{b})(n_{s}+\Delta_{b}^{2})}{n_{b}^{2}+(n_{s}+n_{b})\Delta_{b}^{2}}\\ &\hskip 56.9055pt-\frac{n_{b}^{2}}{\Delta_{b}^{2}}\ln\bigg{(}1+\frac{\Delta_{b}^{2}n_{s}}{n_{b}(n_{b}+\Delta_{b}^{2})}\bigg{)}\bigg{\}}\Bigg{]}^{1/2}\ ,\end{split}$ (24)
@@ 4.2 A signature with significant missing transverse energy
$\begin{split}p\,\,p\ \ \rightarrow&\ \ e^{-}_{7}\,\,e^{+}_{7}+X\ \ \rightarrow\ \ \tau^{-}\,Z\,\tau^{+}\,\,Z+X\ \ \\ &\quad\ \ \rightarrow\tau^{-}\ell^{+}\ell^{-}\,\tau^{-}\ell^{\prime+}\ell^{\prime-}+X\,.\end{split}$ (26)
@@
$n_{s}=53.3\qquad\text{and}\qquad n_{b}=23.8$ (27)
with
$n_{Zh}\simeq 0,\ \ n_{ZZZ}=2.3\ \ \text{and}\ \ n_{t\bar{t}ZZ}=21.5\,,$ (28)
@@ 5 Summary and conclusions
In the absence of direct signals of new physics at colliders, indirect constraints play a crucial role in constructing relevant and motivated BSM theories. This often involves reconciling discrepancies between SM predictions and experimental data, such as those observed in the anomalous magnetic moments of the muon and electron that can be further linked to neutrino physics. Our study exploits precisely this. We examined a simplified extension of the SM to which we added three generations of vector-like leptons, each associated with a specific SM family, and hence determined complementary information compared to previous analyses of VLL effects on the leptonic anomalous magnetic moments. In particular, our work includes the latest experimental constraints on both $\Delta a_{\mu}$ and $\Delta a_{e}$, and it also supplements the analysis of magnetic moments and neutrino masses with an extensive collider investigation of the allowed (favorable) parameter regions.
After exhaustively listing possible representations for VLLs and their interaction Lagrangian, we identified parameter space regions compatible with neutrino data and consistent with measurements of the anomalous magnetic moment of the electron and the muon. We determined that this requires, minimally, the inclusion of two weak VLL singlets (one with the quantum numbers of a right-handed electron and the other with the quantum numbers of a right-handed neutrino), along with a weak VLL doublet with hypercharge $-1/2$. The compatible parameter space regions imply distinct collider signatures, prominently featuring six hard leptons accompanied by either significant total transverse energy or missing transverse energy. We performed a detailed phenomenological analysis to demonstrate the sensitivity of future proton-proton colliders operating at a centre-of-mass energy of 100 TeV to these signatures, focusing on two representative benchmark scenarios with VLL masses in the range of $100-400$ GeV. Our findings underscore that if new physics involves VLLs with properties explaining neutrino data and reconciling discrepancies in anomalous magnetic moments of the light-generation leptons, it will yield signals that will be unmistakable at future proton-proton colliders currently under discussion within the high-energy physics community.